OpenStudy (jiteshmeghwal9):

If 2A=3B=4C, then A:B:C is :-

OpenStudy (jiteshmeghwal9):

@amistre64 @ash2326 plz help

OpenStudy (amistre64):

if 2a = 3b; then what is the ration a:b or rather, a/b ?

OpenStudy (jiteshmeghwal9):

a:b

OpenStudy (amistre64):

what is the ration a:b; solve for a/b

OpenStudy (amistre64):

*ratio

OpenStudy (jiteshmeghwal9):

a/b=3/2

OpenStudy (jiteshmeghwal9):

right ?

OpenStudy (amistre64):

yes; so a=3, b=2 for now, so solve for c knowing that 3b = 4c and b=2

OpenStudy (jiteshmeghwal9):

6/4=c c=3/2

OpenStudy (amistre64):

good now lets look at what we have:\[3:2:\frac{3}2{}\] i dont think they like fractions in these things so would you say we can keep the same ratios by multiplying it all by 2?

OpenStudy (jiteshmeghwal9):

yup

OpenStudy (jiteshmeghwal9):

6:4:3

OpenStudy (amistre64):

i like that result :)

OpenStudy (jiteshmeghwal9):

Thanx !

OpenStudy (amistre64):

lets dbl chk 2A=3B=4C 6 4 3 ----------- 12=12=12 yay!!

OpenStudy (jiteshmeghwal9):

right :)